OpenStudy (anonymous):

solve: 38=2(L)+2(7)... i came up with -25/2L is that correct?

OpenStudy (anonymous):

\[38=2(L)+2(7)\] solve for L?????

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[38=2L+14\] \[2L=38-14\] \[L=24/2\] \[L=12\]

OpenStudy (anonymous):

thanks!